From nch@roe.ac.uk Sat Aug 7 13:59:56 2010 Date: Fri, 6 Aug 2010 16:48:31 +0100 (BST) From: Nigel Hambly To: WFCAM Science Archive Team -- Eckhard Sutorius , Mike Read , Mark Holliman , Nigel Hambly , Nicholas Cross , Rob Blake , Ross Collins Cc: CCs for WSA weekly meeting minutes distribution -- Andrew Lawrence , Jim Emerson , Keith Noddle , Lorenzo Rimoldini , Mike Irwin , Peredur Williams , Bob Mann , Stephen Warren , Tom Kerr Subject: WFAU VDFS Science Archive weekly project meet mins, 6/08/10 Minutes of WFAU VDFS Science Archive meeting: August 6th 2010 ------------------------------------------------------------------------- ------------------------------------------------------------------------- Present: NJC, MAR, NCH, RSC, ETWS Apologies: JPE, KTN, AL, MSH, RGM, RPB DoNM: Friday August 13th in the Vista Hut Actions discharged this week: ----------------------------- ACTION: NJC to check that distortion correction is switched OFF for VIRCam tile in TAN projection in CU13 Actions partly discharged but continuing: ----------------------------------------- ACTION: RSC & NJC to document new operational procedures on the TWiki Continues (nearly done - NJC needs to add some new UDS/SExtractor things) ACTION: NCH to assign actions on the following next time: - Modify entire code base to work with Tiles (Partly discharged - but needs fully testing on the new V1 data). This includes checking code works here on deep stacks, checking quality control/ QBF, ingest, CU7, CU6. Also do we need to do better links between tile detections and pawprint detections or will Provenance etc be good enough. (RSC has added tile qbits; also suggested a preliminary test run of CU7 using shallow tiles first of all). - Change schema / CU4 for new aperture mag names and columns for corrections (partly discharged by ETWS and to be finalised by NJC) - VVV SynopticSource (allowing both SourceXSynopticSourceBestMatch and SourceXDetectionBestMarch) - Updating Filter table with A/E(B-V) and VegaToAB (Need values from STH) Continues Actions carried forward from 30/07/10 meeting --------------------------------------------- ACTION: TBD to draft and MOU on behalf of PSPIs for ESO-SAF deliveries Continues; on hold until at least next PSPI teleconf, and possibly until "Phase 3" workshop at ESO. ACTION: NJC to suggest workshop at next PSPI meeting. ACTION: ETWS to run his auto fitskeys check on the current schema to make sure no vital metadata is missing from the VSA ACTION: RPB to update WFCAMOPENTIME for 09A projects no longer inside their proprietary period. Specific points and new actions: -------------------------------- Project management: Nothing of note this week. WFCAM & VISTA updates: No updates this week. Comments and issues arising from CASU minutes: No new minutes this week. Networking: Another tranche of VIRCam v1.0 products have been made available as of this morning and will be copied up from CASU and ingested immediately. MAR and ETWS noted that one 09B WFCAM night has been identified as having missing detector frames at the JAC end, and this will be looked into today to check on the impact (if any) on UKIDSS DR8. WSA/VSA Operations: ETWS noted -- Finished the implementation of the saturation correction and the new aperMagNoAperCorr values in CU4 and changed the schemas accordingly. -- Moved jpgs and fits files from disk34 to disk36 to free space for general ops and development. -- Minor code changes for WSA monitor pages and fixing the transfer of executable FITS files. Hardware and Systems: Nowt to report this week. Software: RSC noted: "Investigated the append-mode bug in CU7 frame set association and provided a fix to the problem of it leaving out some frames in cases where it shouldn't. Turned on the saturation flagging for deep VISTA stacks as the detections from these catalogues will also be corrected (WFCAM deep data never had this flag applied). Created the detection quality bit flag implementation documentation for the VSA website based on a revised version of the WSA documentation. Tested and fixed a few bugs in NJC's latest developments for programme setup of tiles and also profiled the CU6 best match table creation again following a bug fix that has had the side effect of optimising the last obviously slow part of this code - there are now no more easy gains to be made here." This week NJC has finished off the changes to ProgrammeBuilder to allow it to correctly group pawprints into tiles. NJC has also improved the speed of grouping. NJC has fixed bugs in the updates to the UDS curation and a bug that was introduced into the epoch best-matching. NJC also spotted a problem in the DXS caused by relying on the original deep stack positions supplied by the DXS team rather than using the data as we do with non-surveys and VISTA surveys - i.e. they have observed in new positions. NJC also noted thoughts about DXS colours prompted by the above problem: running ProgrammeSetup, the colours in dxsSource have changed: J-H, H-K rather than J-H, J-K. It was decided to edit the colour part of the automated schema to set it up like the DXS before. This then got NJC thinking about the VHS: VHS is 3 surveys in one, with different filter setups in each, so we may have a similar issue with this survey and need to think about which colours we use: it was decided to put include Y-J, J-H, J-Ks, H-Ks. Regarding aperMagNoAperCor3 etc, it was decided to change extended source colours to use these rather than aperture corrected values. NJC also reported comparing the old CASU tools to the new ones using 24 deep stacks in VIDEO as a test: running imstack, imcore and wcsfit. The data is the same in both cases and generally the average total exposure time is the same. NJC noticed that the noise has significantly decreased, although the seeing seems to have got slightly larger, but overall the number of objects detected is higher. CASU certainly seem to have made some improvements. Survey Data Release: NCH noted that a user has spotted a few missing detector frames in the GCS merging in DR7. Turned out to be an append-mode problem in CU7 which RSC has now fixed; RSC noted that GPS DR6 should be checked for the same problem and notes added to "known issues" for DR6. ACTION: NCH to check GPS DR6 for the odd missing merged frame and make a note on the WSA website as appropriate. After the small hiccup concernig frame set merging, LAS and GCS DR8 are ready to go. ACTION: NCH to do his usual quick science validation/sanity checks on LAS/GCS DR8. We're anticipating delivery of the UDS DR8 stacks later today from Nottingham; the DXS suffered a glitch earlier in the week in that new fields outwith the "required stacks" specification have appeared so NJC is fettling the DXS curation process to handle these, and any subsequent new fields and filters (e.g. if H is ever observed). The GPS survey head Phil Lucas has noted a calibration issue in some Galactic centre fields where there is heavy and highly variable extinction, and some additional quality bit flagging may be required for GPS DR7 et seq. - we await explicit instructions from Phil. Non-survey Data Release: Nothing of note this week. Astrogrid deployment & Data Analysis services: Nothing to report this week. Miscellaneous: That's all for this week. ============================================================= Nigel Hambly Tel: +44-131-668-8234 Institute for Astronomy Fax: +44-131-668-8416 School of Physics and Astronomy University of Edinburgh Email: nch@roe.ac.uk Royal Observatory Blackford Hill Edinburgh EH9 3HJ The University of Edinburgh is a charitable body, registered in Scotland, with registration number SC005336. =============================================================